OpenStudy (anonymous):

Solve: -4(x + 1) - 3 = -3(x - 4) A.) -19 B.) 19 C.) 2 D.) 0

OpenStudy (anonymous):

A.) -19

OpenStudy (anonymous):

Do you need me to show you how I got the answer?

OpenStudy (anonymous):

Yes please

OpenStudy (anonymous):

-4(x+1) -3 = -3(x-4) -4x -4 -3 = -3x + 12 -4x -7 = -3x + 12 subtract 12 from both sides -4x -19 = -3x add 4x to each side -19 = x Does that make sense?

OpenStudy (anonymous):

thank you